# Observational constraints on degeneracy of non-canonical inflation driven by quartic potential

### v2

@@ Abstract
Here, the quartic inflationary potential $V(\phi)=\frac{\lambda}{4}\phi^{4}$ in non-canonical framework with a power-law Lagrangian is investigated. We demonstrate that the predictions of this potential in non-canonical framework align with the observational data of Planck 2018. We determine how the predictions of the model vary with changes in the non-canonical parameter $\alpha$ and the number of $e$-folds $N$. The sound speed, non-Gaussianity parameter, scalar spectral index, and tensor-to-scalar ratio are all influenced by $\alpha$. However, the scalar spectral index exhibits degeneracy with respect to changes in $\alpha$. It is found that, this degeneracy does not break through reheating consideration. But the length of the reheating epoch constrains the length of the inflation period and the value of $\alpha$. Therefore, we investigate relic gravitational waves (GWs) and show that, the degeneracy of the model with respect to the $\alpha$ parameter can be broken using the relic GWs. Additionally, the density of relic GWs falls within the sensitivity range of GWs detectors for the specific $e$-folds number between 55 and 55.7.
@@ I Introduction
The Big Bang theory of cosmology provides an explanation for structure and evolution of the universe. The theory of cosmic inflation was proposed to complement the Big Bang theory. Inflation is an unknown phenomenon during which the universe expands exponentially. In Einstein’s gravitation framework, a negative pressure is required to justify this accelerated expansion. The quantum fields can be the source of negative pressure. The scalar field associated with inflation is usually called inflaton. The nature of the inflaton is still unknown. However, the exact causes of cosmic inflation have remained unknown. It has been suggested that the discovery of certain particles in fundamental particle physics can explain the cause of cosmic inflation Starobinsky:1980 ; Guth:1981 ; Linde:1982 ; Albrecht:1982 ; Linde:1990 ; Baumann:2009 . One of the predictions of inflationary cosmology is the generation of primordial scalar fluctuations. The growth of these fluctuations can lead to the formation of large-scale structures. Additionally, the perturbations caused by scalar and tensor fluctuations during inflation can impact the cosmic microwave background (CMB) radiation. These effects are utilized to determine the scalar spectral index ($n_{s}$) and the tensor-to-scalar ratio ($r$), which are two important observational parameters Liddle:2000 ; Baumann:2009 . These parameters are commonly used to test the validity of inflationary models. However, similar values of $n_{s}$ or $r$ can be obtained by adjusting one or more free parameters in different models Liddle:2000 ; BICEP2:2018 . This leads to a problem of indistinguishability among the models. Therefore, relying solely on observations of CMB radiation is insufficient to thoroughly investigate and differentiate between various inflationary models Kallosh:2013a ; Kallosh:2013 ; akrami:2020 . Other observable parameters, such as the parameters of reheating epoch and relic gravitational waves (GWs), should be considered to effectively discern these models Grishchuk:1974 ; Starobinsky:1979 ; Starobinsky:1981 .
@@
The universe has undergone various periods following inflation, each characterized by a specific equation of state parameter, including radiation-dominated and matter-dominated epochs. After the end of inflation and prior to the onset of the radiation-dominated period, the universe enters a reheating phase. Despite significant theoretical advancements in understanding the early universe, the physics of the reheating epoch remains poorly understood. At the end of inflation, the kinetic energy of the inflaton overcomes its potential energy. The inflaton then begins to oscillate around its potential minimum, gradually giving up its energy to other fields. However, if this process is slow, the reheating takes place perturbatively and the scalar field oscillates for a much longer time and gradually releases its energy in the form of radiation in the universe Baumann:2009 ; Kofman:1994 ; Kofman:1996 ; Gialamas:2020 ; Odintsov:2023 . In this scenario, the scalar field oscillates for an extended period, and the equation of state parameter of this oscillatory phase is determined by the form of the inflaton potential at its minimum value Baumann:2009 . Considering the reheating epoch can also be useful for verifying inflationary models. Descriptive parameters of the reheating period, such as the reheating duration ($N_{re}$), reheating temperature ($T_{re}$), and the equation of state parameter ($\omega_{re}$), are usually analyzed for this epoch.
@@
In 2010, Martin and Ringeval Martin:2010 examined the constraints of CMB radiation on the reheating temperature. In 2014, Dai et al. LiangDai:2014 investigated observational constraints on the inflationary power-law potentials and found that, the reheating constraints could impose more restrictions on this class of potentials too. In 2016, Eshaghi et al. Eshaghi:2016 studied the $\alpha$-attractors inflationary models, including T-model and E-model. They calculated the predictions of these models for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$, and then applied limits based on the observational data of Planck 2015. Furthermore, they determined a range for $\alpha$ based on the reheating constraints. In 2021, Mishra et al. Mishra:2021 examined the parameters of the reheating epoch and relic gravitational waves for the same two types of $\alpha$-attractors inflationary models and tried to break the degeneracies of the predictions of the model with respect to the model parameters.
@@
In this work, we are interested in studying the reheating and relic GWs constraints on the quartic inflationary potential, $V(\phi)=\frac{\lambda}{4}\phi^{4}$, in non-canonical framework with power-law Lagrangian. To do so, in section II a brief formulation of the inflationary model with quartic potential in the non-canonical framework is explained. In section III, the parameters of the reheating epoch for the model are examined. In section IV, the density of relic GWs predicted by the model is compared with the sensitivity curves of GWs detectors. Finally, in section V, the conclusions of our work are presented.
@@ II Non-canonical inflation
where $g$ is determinant of the metric tensor $g_{\mu\nu}$ and $R$ is the Ricci scalar. Also the Lagrangian density $\mathcal{L}(X,\phi)$ is a function of the scalar field $\phi$ and the kinetic energy $X\equiv\frac{1}{2}g_{\mu\nu}\partial^{\mu}\phi\partial^{\nu}\phi$. In this work, we consider the power-law form of the Lagrangian density as follows Panotopoulos-2007 ; Unnikrishnan:2012 ; Rezazadeh:2015 ; Mishra:2022 ; Heydari:2024a ; Heydari:2024b
@@
$\rho_{\phi}=\left(\frac{\partial{\cal L}}{\partial X}\right)\left(2X\right)-{\cal L}=\left({2\alpha-1}\right)X{\left({\frac{X}{{{M^{4}}}}}\right)^{\alpha-1}}+V(\phi),$ (3)
${p_{\phi}}={\cal L}=X{\left({\frac{X}{{{M^{4}}}}}\right)^{\alpha-1}}-V(\phi).$ (4)
@@
$\epsilon\equiv-\frac{\dot{H}}{H^{2}}~{},~{}\eta\equiv\epsilon-\frac{\dot{\epsilon}}{2H\epsilon}.$ (9)
@@
$\epsilon\simeq\epsilon^{\rm NC}_{V}\equiv\left[\frac{1}{\alpha}\left(\frac{3M^{4}}{V\left(\phi\right)}\right)^{\alpha-1}\left(\frac{M_{p}{V}^{\prime}\left(\phi\right)}{\sqrt{2}V\left(\phi\right)}\right)^{2\alpha}\right]^{\frac{1}{2\alpha-1}},$ (12)
$\eta\simeq\eta^{\rm NC}_{V}\equiv\left(\frac{\alpha\epsilon}{2\alpha-1}\right)\left(2\frac{V\left(\phi\right){V}^{\prime\prime}\left(\phi\right)}{{V}^{\prime}\left(\phi\right)^{2}}-1\right),$ (13)
where the superscript ”NC” denotes the non-canonical framework. It is notable that, for $\alpha=1$, Eqs. (12) and (13) reduce to $\epsilon_{V}=\frac{M_{p}^{2}}{2}\left(\frac{V^{\prime}}{V}\right)^{2}$ and $\eta_{V}=M_{p}^{2}\left(\frac{V^{\prime\prime}}{V}\right)$ which describe the canonical form of the potential slow-roll parameters.
As for the dynamics of scalar and tensor perturbations under the slow-roll approximations within the non-canonical framework, following Garriga:1999 , the power spectrum of scalar perturbation is given by
@@
In order to prevent the classical and phantom instabilities, it is necessary that $0\leq c_{s}^{2}\leq 1$ and consequently from Eq. (16) we have $\alpha\geq 1$. In the following, the scalar spectral index $n_{s}$ is obtained from the scalar power spectrum and in the non-canonical framework, it can be written in terms of slow-roll parameters as follows
@@
The value of the scalar spectral index $n_{s}$ is constrained by Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO at 68% and 95% CL as bk18 ; Paoletti:2022 ; akrami:2018
@@
${\cal P}_{t}=\frac{2H^{2}}{\pi^{2}M_{p}^{2}}\Big{|}_{k=aH}.$ (19)
Using the power spectrum of tensor perturbations, the tensor spectral index is calculated in terms of first slow-roll parameter as
@@
In what follows, we review the analytical calculations of Unnikrishnan:2012 for power-law potential in the non-canonical framework containing the power-law Lagrangian (2). The power-law potential is given by
@@
The inflationary $e$-fold number from the end of inflation to the CMB horizon crossing moment is given by
@@
According to Eq. (17) and using Eqs. (28) and (29), the scalar spectral index for the power-law potential in non-canonical framework with power-law Lagrangian reads
@@
Thereafter, inserting Eq. (28) in (31) and using Eq. (20), the tensor spectral index for power-law potential is calculated as follows
@@
where $\lambda=0.1$ is dimensionless self-interaction parameter Tanabashi:2018 ; Workman:2022 . Note that our motivation for considering the potential (35) comes back to this fact that the prediction of this potential in the standard model is completely ruled out by the Planck observations. But in what follows, we will show that within the framework of non-canonical inflation for the special ranges of $\alpha$ parameter, the prediction of quartic potential can lie inside the allowed regions of the Planck 2018 data (see the $r-n_{s}$ plane in Fig. 3). For the quartic potential (35), $V_{0}=\lambda/4$ and $n=4$, the scalar spectral index (30) is simplified as follows
@@
It should be noted that, if $\alpha=1$, all above equations return to the standard inflation.
@@
In Fig. 1, using Eq. (36), the scalar spectral index $n_{s}$ is plotted as a function of the non-canonical parameter $\alpha$ for various values of the inflationary $e$-fold number. The permissible range for the scalar spectral index $n_{s}$ according to Planck 2018 data in 68% (95%) CL is depicted in dark (light) blue in the background. It can be seen that the scalar spectral index parameter in this model for $N=50$ (solid red curve) cannot be consistent with the Planck observational data. Our estimations indicate that for $N=55$, the prediction of the model falls within the 95% CL for $\alpha\geq 5$ (see the dashed blue line in Fig. 1). Note that, as the value of $\alpha$ increases, the value of $n_{s}$ remains almost constant, as depicted in Fig. 1. If we set $N=60$ (dashed red line), the estimated value for $n_{s}$ can be consistent with the 68% and 95% CL when $\alpha\geq 2$ and $\alpha\geq 10$, respectively. Similar to previous case, the degeneracy of the scalar spectral index with respect to $\alpha$ parameter occurs at higher values of $\alpha$. For a larger $e$-fold number, our model aligns with observations at smaller $\alpha$ values. However, the degeneracy of model persists at larger $\alpha$ values. So making a correct analysis solely based on CMB observations is impossible.
@@
Figure 1: The scalar spectral index $n_{s}$ versus non-canonical parameter $\alpha$ for different $e$-folds number $N=50$ (solid red curve), $N=55$ (dashed blue curve), $N=60$ (dashed red curve), $N=65$ (dashed black curve), and $N=70$ (solid brown curve). The dark and light blue regions show, respectively, the 68% and 95% CL of Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO data.
@@
In Fig. 2, using Eq. (37) the tensor-to-scalar ratio $r$ is depicted in terms of $\alpha$ parameter for various duration of inflationary era. Furthermore, the allowed area for $r$ at 95% CL according to the recent data of Planck’s measurements (22) is marked in light blue in this figure. As depicted in Fig. 2, for $N=70$, the prediction of the model falls within the allowed region for $\alpha\geq 10$. Additionally, the cases associated with smaller $N$ can satisfy observational constraints at higher values of $\alpha$, as it can be seen from Fig. 2.
@@
Figure 2: The tensor-to-scalar ratio $r$ versus the non-canonical parameter $\alpha$ for different $e$-folds number $N$. The light blue region shows $r\leq 0.036$ according to Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO at 95% CL.
The allowed ranges for $\alpha$ parameter in $r-n_{s}$ diagram in light of Planck 2018 data for different values of $N$ are shown in Fig. 3 and Table 1. According to these results, for $N=60$, the $\alpha$ parameter must be greater than 18.8 to be in the allowed area. The $f_{\rm NL}^{\rm equil}$ parameter also limits the maximum value of $\alpha$ to 130 in this model (see Table 1).
@@
Figure 3: The tensor-to-scalar ratio $r$ versus the scalar spectral index $n_{s}$ with varying $\alpha$ parameter and for different $e$-folds number $N=$ 55 (solid red curve), 55.3 (solid black curve), 55.7 (solid yellow curve), 60 (dashed red curve), 65 (dashed black curve), and 70 (brown curve). Dark (light) green area represents the 68% (95%) CL of Planck 2018 TT, TE, EE+low E+lensing and dark (light) blue one shows the 68% (95%) CL of Planck 2018 TT, TE, EE+low E+lensing+BK18+BAO in the background.
In Table 1, the computed allowed ranges for the non-canonical parameters $\alpha$ and $\mu$, using the constraints of $c_{s}^{2}+r+n_{s}+f_{\rm NL}^{\rm equil}$ for different $e$-folds number $N$ are expressed. For instance, if the duration of the inflationary era is assumed to be 55 $e$-folds, the $r-n_{s}$ curve is placed within the allowed range of Planck 2018 data for $\alpha\geq 110$. In this calculation, the dimensionless coefficient of self-interaction $\lambda=0.1$ is considered.
@@
Table 1: The allowed ranges for the non-canonical parameters of the model ($\alpha$, $\mu$) according to the constraints of $c_{s}^{2}+r+n_{s}+f_{\rm NL}^{\rm equil}$ for different $e$-folds number $N$.
  $N$   $\alpha$   $\mu~{}(\times 10^{-5})$
  $55$   $110\leq\alpha\leq 130$   $2.23\leq\mu\leq 2.30$
  $60$   $18.8\leq\alpha\leq 130$   $2.07\leq\mu\leq 2.59$
  $65$   $11.3\leq\alpha\leq 130$   $1.93\leq\mu\leq 2.34$
  $70$   $8.6\leq\alpha\leq 130$   $1.812\leq\mu\leq 2.06$

### v3

# Breaking the degeneracy of non-canonical quartic inflation by reheating considerations

@@ Abstract
Here, the quartic inflationary potential $V(\phi)=\frac{\lambda}{4}\phi^{4}$ within a non-canonical framework characterized by a power-law Lagrangian is investigated. We demonstrate that the predictions of this model align with the Planck 2018 observational data. We explore how the predictions of the model depend on the non-canonical parameter $\alpha$ and the number of $e$-folds $N$. Notably, the sound speed, non-Gaussianity parameter, scalar spectral index, and tensor-to-scalar ratio are all affected by variations in $\alpha$. However, the scalar spectral index exhibits a degeneracy with respect to variation in $\alpha$, which can be broken by incorporating reheating consideration. By applying a combination of theoretical and observational constraints on $(r-n_{\rm s})$, non-Gaussianity, and reheating parameters, we find that the duration of inflation is constrained to the range $55\leq N\leq 55.7$ $e$-folds for $60\leq\alpha\leq 130$. Finally, we investigate relic gravitational waves and demonstrate that their energy density spectrum falls within the sensitivity range of gravitational waves detectors for this constrained range of $e$-folds.
@@ I Introduction
The Big Bang theory of cosmology provides an explanation for structure and evolution of the universe. The theory of cosmic inflation was proposed to complement the Big Bang theory. Inflation is an unknown phenomenon during which the universe expands exponentially. In Einstein’s gravitation framework, a negative pressure is required to justify this accelerated expansion. The quantum fields can be the source of negative pressure. The scalar field associated with inflation is usually called inflaton. The nature of the inflaton is still unknown. However, the exact causes of cosmic inflation have remained unknown. It has been suggested that the discovery of certain particles in fundamental particle physics can explain the cause of cosmic inflation Starobinsky:1980 ; Guth:1981 ; Linde:1982 ; Albrecht:1982 ; Linde:1990 ; Baumann:2009 . One of the predictions of inflationary cosmology is the generation of primordial scalar fluctuations. The growth of these fluctuations can lead to the formation of large scale structures. Additionally, the perturbations caused by scalar and tensor fluctuations during inflation can impact the cosmic microwave background (CMB) radiation. These effects are utilized to determine the scalar spectral index ($n_{\rm s}$) and the tensor-to-scalar ratio ($r$), which are two important observational parameters of inflationLiddle:2000 ; Baumann:2009 . These parameters are commonly used to test the validity of inflationary models. However, similar values of $n_{\rm s}$ or $r$ can be obtained by adjusting one or more free parameters in different models Liddle:2000 ; BICEP2:2018 . This leads to a problem of indistinguishability among the models. Therefore, relying solely on observations of CMB radiation is insufficient to thoroughly investigate and differentiate between various inflationary models Kallosh:2013a ; Kallosh:2013 ; akrami:2020 . Other observable parameters, such as the parameters of reheating epoch and relic gravitational waves (GWs), should be considered to effectively discern these models Grishchuk:1974 ; Starobinsky:1979 ; Starobinsky:1981 .
@@
The universe has undergone various periods following inflation, each characterized by a specific equation of state parameter, including radiation-dominated and matter-dominated epochs. After the end of inflation and prior to the onset of the radiation-dominated period, the universe enters a reheating phase. Despite significant theoretical advancements in understanding the early universe, the physics of the reheating epoch remains poorly understood. At the end of inflation, the kinetic energy of the inflaton overcomes its potential energy. The inflaton then begins to oscillate around its potential minimum, gradually giving up its energy to other fields. However, if this process is slow, the reheating takes place perturbatively and the scalar field oscillates for a much longer time and gradually releases its energy in the form of radiation in the universe Baumann:2009 ; Kofman:1994 ; Kofman:1996 ; Gialamas:2020 ; Odintsov:2023 . In this scenario, the scalar field oscillates for an extended period, and the equation of state parameter of this oscillatory phase is determined by the form of the inflaton potential at its minimum value Baumann:2009 . Considering the reheating epoch can also be useful for verifying inflationary models. Descriptive parameters of the reheating period, such as the reheating duration ($N_{\rm re}$), reheating temperature ($T_{\rm re}$), and the equation of state parameter ($\omega_{\rm re}$), are usually analyzed for this epoch.
@@
In Martin:2010 , Martin and Ringeval examined the constraints of CMB radiation on the reheating temperature. In LiangDai:2014 , Dai et al. investigated observational constraints on the inflationary power-law potentials and found that, the reheating constraints could impose more restrictions on this class of potentials too. In Eshaghi:2016 , Eshaghi et al. studied the $\alpha$-attractors inflationary models, including T-model and E-model. They calculated the predictions of these models for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$, and then applied limits based on the observational data of Planck 2015. Furthermore, they determined a range for $\alpha$ based on the reheating constraints. In Mishra:2021 , Mishra et al. showed that reheating parameters, along with the spectrum of relic GWs, can be used to distinguish between canonical and non-canonical inflation, and to break the degeneracies between two $\alpha$-attractors models with respect to their free parameters.
@@
All mentioned in above motivate us to study the observational constraints on degeneracy of the quartic inflationary potential, $V(\phi)=\frac{\lambda}{4}\phi^{4}$, in non-canonical framework with power-law Lagrangian. To do so, in section II a brief formulation of the inflationary model with quartic potential in the non-canonical framework is explained. In section III, the parameters of the reheating epoch for the model are examined. In section IV, the density of relic GWs predicted by the model is compared with the sensitivity curves of GWs detectors. Finally, in section V, the conclusions of our work are presented.
@@ II Non-canonical inflation
where $g$ is determinant of the metric tensor $g_{\mu\nu}$ and $R$ is the Ricci scalar. Also the Lagrangian density $\mathcal{L}(X,\phi)$ is a function of the kinetic energy $X\equiv\frac{1}{2}g_{\mu\nu}\partial^{\mu}\phi\partial^{\nu}\phi$ and the scalar field $\phi$. In this work, we consider the power-law form of the Lagrangian density as follows Panotopoulos-2007 ; Unnikrishnan:2012 ; Rezazadeh:2015 ; Mishra:2022 ; Heydari:2024a ; Heydari:2024b ; Heydari:2024c
@@
$\rho_{\phi}\equiv\left(\frac{\partial{\cal L}}{\partial X}\right)\left(2X\right)-{\cal L}=\left({2\alpha-1}\right)X{\left({\frac{X}{{{M^{4}}}}}\right)^{\alpha-1}}+V(\phi),$ (3)
${p_{\phi}}\equiv{\cal L}=X{\left({\frac{X}{{{M^{4}}}}}\right)^{\alpha-1}}-V(\phi).$ (4)
@@
$\epsilon\equiv-\frac{\dot{H}}{H^{2}}~{},~{}\eta\equiv-\frac{\ddot{\phi}}{H\dot{\phi}}=\epsilon-\frac{\dot{\epsilon}}{2H\epsilon}.$ (9)
@@
$\epsilon\simeq\left[\frac{1}{\alpha}\left(\frac{3M^{4}}{V\left(\phi\right)}\right)^{\alpha-1}\left(\frac{M_{p}{V}^{\prime}\left(\phi\right)}{\sqrt{2}V\left(\phi\right)}\right)^{2\alpha}\right]^{\frac{1}{2\alpha-1}},$ (12)
$\eta\simeq\left(\frac{\alpha\epsilon}{2\alpha-1}\right)\left(2\frac{V\left(\phi\right){V}^{\prime\prime}\left(\phi\right)}{{V}^{\prime}\left(\phi\right)^{2}}-1\right).$ (13)
It is notable that, for $\alpha=1$, Eqs. (12) and (13) reduce to $\epsilon\simeq\epsilon_{V}$ and $\eta\simeq\eta_{V}-\epsilon_{V}$, where $\epsilon_{V}\equiv\frac{M_{p}^{2}}{2}\left(\frac{V^{\prime}}{V}\right)^{2}$ and $\eta_{V}\equiv M_{p}^{2}\left(\frac{V^{\prime\prime}}{V}\right)$ represent the canonical forms of the first and second potential slow-roll parameters, respectively.
As for the dynamics of scalar and tensor perturbations under the slow-roll approximation within the non-canonical framework, following Garriga:1999 , the power spectrum of scalar perturbation is given by
@@
In order to prevent the classical and phantom instabilities, it is necessary that $0\leq c_{s}^{2}\leq 1$ and consequently from Eq. (16) we have $\alpha\geq 1$. In the following, the scalar spectral index $n_{s}$ is obtained from the scalar power spectrum in the non-canonical framework and it can be written in terms of slow-roll parameters as follows
@@
The value of the scalar spectral index $n_{\rm s}$ is constrained by Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO at 68% and 95% CL as bk18 ; Paoletti:2022 ; akrami:2018
@@
${\cal P}_{t}=\frac{2H^{2}}{\pi^{2}M_{p}^{2}}\Big{|}_{k=aH},$ (19)
and the tensor spectral index is calculated in terms of first slow-roll parameter as
@@
In what follows, we consider the power-law potential given by
@@
The inflationary $e$-folds number from the end of inflation to the CMB horizon crossing moment is given by
@@
Replacing Eq. (28) into (12), (13) and using Eq. (17), the scalar spectral index for the power-law potential (24) in non-canonical framework with power-law Lagrangian reads
@@
Thereafter, inserting Eq. (28) in (12) and using Eq. (20), the tensor spectral index for the power-law potential (24) is calculated as follows
@@
where $\lambda=0.1$ is dimensionless self-interaction parameter Tanabashi:2018 ; Workman:2022 . For the quartic potential (35), with the help of Eqs. (10) and (28), in Fig. 1 we plot the variations of the background quantities including the scalar field $\phi$ and the Hubble parameter $H$ as functions of the $e$-folds number $N$ for different values of $\alpha$. The figure shows that both $\phi$ and $H$ decrease during inflation, and that increasing $\alpha$ results in lower values of both quantities at a given $N$. This behaviour indicates that larger values of $\alpha$ correspond to a lower inflationary energy scale. It should be noted that the end of inflation is defined at $N=0$.
Figure 1: Evolutions of (a) the scalar field $\phi$ and (b) and the Hubble parameter $H$, for different values of $\alpha$. The green, blue, black dashed, and red lines represent $\alpha=11.3,18.8,110$ and $130$, respectively. The end of inflation is set at $N=0$.
Note that our motivation for considering the potential (35) comes back to this fact that the prediction of this potential in the standard model is completely ruled out by the Planck observations bk18 ; Paoletti:2022 ; akrami:2018 . But in what follows, we will show that within the framework of non-canonical inflation for the special ranges of $\alpha$ parameter, the prediction of quartic potential can lie inside the allowed regions of the Planck 2018 data (see the $r-n_{\rm s}$ panel in Fig. 4). For the quartic potential (35), $V_{0}=\lambda/4$ and $n=4$, the scalar spectral index (30) is simplified as follows
@@
In Fig. 2, using Eq. (36), the scalar spectral index $n_{\rm s}$ is plotted as a function of the non-canonical parameter $\alpha$ for various values of the inflationary $e$-folds number. The permissible range for the scalar spectral index $n_{\rm s}$ according to Planck 2018 data in 68% (95%) CL is depicted in dark (light) blue in the background. Our estimations indicate that for $N=55$, the prediction of the model falls within the 95% CL when $\alpha\geq 5$ (see the solid red line in Fig. 2). Note that, as the value of $\alpha$ increases, the value of $n_{\rm s}$ remains almost constant, as depicted in Fig. 2. If we set $N=60$ (dashed black line), the estimated value for $n_{\rm s}$ can be consistent with the 68% and 95% CL when $\alpha\geq 2$ and $\alpha\geq 10$, respectively. Similar to previous case, the degeneracy of the scalar spectral index with respect to $\alpha$ parameter occurs at higher values of $\alpha$. For a larger $e$-folds number, our model aligns with observations at smaller $\alpha$ values. However, the degeneracy of model persists at larger $\alpha$ values. So making a correct analysis solely based on CMB observations is impossible.
@@
Figure 2: The scalar spectral index $n_{\rm s}$ versus non-canonical parameter $\alpha$ for different $e$-folds number $N=55$ (solid red curve), $N=60$ (dashed black curve), $N=65$ (dashed red curve). The dark and light blue regions show, respectively, the 68% and 95% CL of Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO data bk18 ; Paoletti:2022 ; akrami:2018 .
@@
In Fig. 3, using Eq. (37) the tensor-to-scalar ratio $r$ is depicted in terms of $\alpha$ parameter for various duration of inflationary era. Furthermore, the allowed area for $r$ at 95% CL according to the recent data of Planck’s measurements (22) is marked in light blue in this figure. As depicted in Fig. 3, for $N=65$, the prediction of the model falls within the allowed region for $\alpha\geq 10$. Additionally, the cases associated with smaller $N$ can satisfy observational constraints at higher values of $\alpha$, as it can be seen from Fig. 3.
@@
Figure 3: The tensor-to-scalar ratio $r$ versus the non-canonical parameter $\alpha$ for different $e$-folds number $N$. The light blue region shows $r\leq 0.036$ according to Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO at 95% CL bk18 ; Paoletti:2022 ; akrami:2018 .
Figure 4 shows the $r-n_{\rm s}$ diagram in light of Planck and BICEP/Keck 2018 data (TT, TE, EE+lowE+lensing+BK18 +BAO) for different values of $N$ and varying $\alpha$. Note that on each curve, the parameter $\alpha$ increases from top to bottom.
@@
Figure 4: The tensor-to-scalar ratio $r$ versus the scalar spectral index $n_{\rm s}$ with varying $\alpha$ parameter and for different $e$-folds number $N=$ 55 (solid red curve), 60 (dashed black curve), and 65 (dashed red curve). Dark (light) green area represents the 68% (95%) CL of Planck 2018 TT, TE, EE+low E+lensing and dark (light) blue one shows the 68% (95%) CL of Planck 2018 TT, TE, EE+low E+lensing+BK18+BAO in the background.
@@
In Table 1, the computed allowed ranges for the non-canonical parameters $\alpha$ and $\mu$, using the constraints of $(r-n_{s})+f_{\rm NL}^{\rm equil}$ for different $e$-folds number $N$ are expressed. For instance, for the $e$-folds number $N=55$, the allowed range for the $\alpha$ parameter is $110\leq\alpha\leq 130$ at 95% CL. Note that the $f_{\rm NL}^{\rm equil}$ parameter limits the maximum value of $\alpha$ to 130 in this model.
Table 1: The allowed ranges for the model parameters $\alpha$ and $\mu\equiv M/M_{p}$ related to the permitted $e$-folds number $N$ in the non-canonical quartic inflationary model, considering a combination of various constraints from Planck and BICEP/Keck 2018 data on ($r-n_{\rm s}$), the equilateral non-Gaussianity parameter $f_{\rm NL}^{\rm equil}$, reheating $(\omega_{\rm re}+N_{\rm re}+T_{\rm re}$), and relic GWs. Here, the parameter $\mu$ is estimated from Eq. (38).
$N$ $(r-n_{\rm s})+f_{\rm NL}^{\rm equil}$ $(r-n_{\rm s})+f_{\rm NL}^{\rm equil}$ $+\omega_{\rm re}+N_{\rm re}+T_{\rm re}$
$(r-n_{\rm s})+f_{\rm NL}^{\rm equil}$ $+\omega_{\rm re}+N_{\rm re}+T_{\rm re}+{\rm GWs}$
$\alpha(95\%\text{ CL})$ $\alpha(68\%\text{ CL})$ $\mu(\times 10^{-5})~{}(95\%\text{ CL})$ $\mu(\times 10^{-5})~{}(68\%\text{ CL})$ $\alpha(95\%\text{ CL})$ $\mu(\times 10^{-5})~{}(95\%\text{ CL})$
$55$ $[110,130]$- $[2.23,2.30]$- $[110,130]$ $[2.23,2.30]$
$55.3$ $[80,130]$- $[2.22,2.40]$- $[80,130]$ $[2.22,2.40]$
$55.7$ $[60,130]$- $[2.21,2.49]$- $[60,65]$ $[2.46,2.49]$
$60$ $[18.8,130]$- $[2.07,2.59]$- -- --
$65$ $[11.3,130]$ $[25,130]$ $[1.93,2.34]$ $[1.93,2.39]$- --